# Global Prediction of COVID-19 Variant Emergence Using Dynamics-Informed Graph Neural Networks

### v1

@@ I Introduction
The COVID-19 pandemic presented an unprecedented global health crisis, severely affecting millions of people worldwide and demanding swift and effective responses from governments and healthcare providers [1]. As the pandemic progressed, multiple variants of COVID-19 emerged, each possessing unique genetic mutations that can significantly impact transmissibility, virulence, and even vaccine efficacy. During much of the COVID-19 epidemic, the surge in cases and severe outcomes (hospitalizations and deaths) have been driven by the emergence of new variants [2]. Consequently, monitoring the emergence and spread of these variants is crucial for devising appropriate public health measures and optimizing containment strategies [3]. A critical factor driving the surge in a given region is the time of arrival of the new variant [4]. We can observe how a new variant, that has not appeared in region $A$ yet, spreads in region $B$. We can study the spread in the region $B$ to understand the properties of this new variant. If it is a highly transmissible or immune-evading variant [5], we can expect it to spread in the region $A$ eventually. However, when it happens in region $A$ precisely remains unknown. Therefore, making a good prediction of arrival time would lead to a more effective preparation and resource management.
@@
It naturally follows from the above description that Problem 1 is a classification problem while Problem 2 is a regression problem. In practical scenarios, we envision addressing Problem 1 first, and only tackling Problem 2 if the outcome is positive. This is also important for the sake of evaluation as we can answer Problem 2 only if the variant actually arrives in the region of interest. Existing work has focused on the problem of predicting the prevalence of a variant when it has already arrived in the region of interest [6, 7, 8]. Specifically, logistic regression turns out to provide a good estimate of the prevalence over time [9, 10]. However, these techniques require the variant to have a non-zero proportion already in the region of interest. To the best of our knowledge, no prior work exists predicting the delay to reach a certain prevalence even when there is zero prevalence in the region. We show that Problem 1 (Emergence prediction) is a simpler problem (albeit, non-trivial) that can be solved by a decision tree. However, Problem 2 (delay detection) is more difficult to solve and requires more sophisticated techniques.
When a new variant emerges with evolutionary favorable properties, the only way it can be transferred to a different region is through a host (human). This encourages the idea of using an underlying network of mobility to address the proposed problems. Since reaching a certain prevalence may depend on the other currently circulating variants, their dynamics (how fast one variant can spread over others) also play a role. Therefore, we propose a variant dynamics-informed Graph Neural Network (GNN) that utilizes a network of mobility and features inspired by variant dynamics to solve the proposed problems. This is different than typical Physics-Informed Neural Networks (PINNs) where the dynamics act as a regularization for the loss [11]. Here, we show that our approach of constructing appropriate features results in lower errors than incorporating the dynamics in the loss function.
Our key contributions represent a novel effort in addressing the challenge of predicting the emergence of COVID-19 variants at a global level, and in doing so establishing a new benchmark for evaluating Emergence Detection and Delay Prediction on the CoVariants dataset [12]. More specifically, our contributions are as follows:
@@ II-A Variants and Variant Dynamics
In the context of infectious diseases, the term ”variant” refers to a version of a virus with some changes in its genetic material, known as genome [13]. These changes happen through genetic mutations and can affect the virus’s characteristics, like how easily it spreads, how severe the illness it causes is, and whether it can evade the immune system or not. Variants can be categorized in multiple ways based on genetic differences, and the significance of these differences can vary. Some categorizations focus on a common ancestor, while others may specifically highlight mutations in key regions of the virus’s genome [13]. In our study, we rely on the categorizations provided by the CoVariants dataset, discussed further in Section IV-A.
@@
Furthermore, it’s worth noting that the proportions of different variants can be analyzed over time. Some studies have observed that these proportions follow a straight line when plotted on a semi-logarithmic scale. Therefore, they leverage this fact to understand the dynamics of variant emergence and prevalence in populations [14]. We also observe this in our data as shown in Figure 2. However, to the best of our knowledge, existing studies do not derive any dynamics capable of predicting delays in the emergence of variants between multiple regions/countries.
@@ II-B Graph Neural Networks and Epidemics
In response to the COVID-19 pandemic, many efforts were made to develop forecasting models capable of predicting the progression and future trends of the COVID-19 pandemic [15, 16, 17]. However, these forecasting approaches have primarily focused on predicting overall case counts, hospitalizations, or mortality rates, rather than predicting delays in variant emergence. The ability to predict the timing of variant-specific outbreaks can significantly enhance public health preparedness and response strategies, allowing for timely interventions such as targeted vaccination campaigns and tailored public health measures [3].
Many Deep Learning (DL) techniques have been explored for predicting the dynamics of infectious diseases. While many models use methods like GNNs [18], and PINNs [19] for forecasting, some approaches have also focused on genetic sequence modeling for variants [20, 21].
E. Kharazmi et al. [22], use PINNs in epidemiological models to better capture the dynamics through the underlying Ordinary Differential Equations (ODEs). These works are normally variations of the classical Susceptible-Infectious-Removed (SIR) model with additional compartments [23]. Extending this, A. Rodríguez et al. [24] propose a framework for forecasting, Epidemiologically-informed Neural Networks (EINNs) that integrates epidemiological knowledge embedded in ODE models into neural networks. The core idea is utilizing the learned representations from a PINN using transfer learning along with a neural network that ingests multi-modal data. Other extensions include Epi-DNNs from X. Ning et al. [25], in which the neural network expresses the unknown parameters in the compartmental model, and the ODEs are solved using the Runge-Kutta method.
Since infectious diseases inherently present across multiple regions over a period of time, network models, like GNNs help in capturing the low-level graph effects. Works [26, 27, 28, 29, 30], including Gao et al. [27] rely on GNNs to predict the total number of cases for a fixed period in the future using a spatio-temporal attention network. These models typically map locations (e.g., a county or a state) to nodes on a graph and construct the edges based on geographical proximity and demographic similarity between locations. As an extension, [31, 32] combines spatio-temporal GNNs with PINNs. [33, 34] use agent-based models on top of these to further capture low-level interactions between populations in different compartments.
While some works do focus on using the spatio-temporal networks with location aware features [35], including mobility data [36], to the best of our knowledge, no studies have attempted to predict the emergence of a variant in one country based on its occurrence in another (Problems 1 & 2), through variant informed dynamics.
@@ III-B Variant Growth Rates
We first derive the dynamics of two competing variants in one region. While logistic regression has been widely used to fit these dynamics, for completeness, we present a derivation. We start by using a model of infectious diseases that generalizes a large class of models [37]:
@@
The above further assumes that the mean serial intervals of the two variants are close, i.e., $b_{i}=b_{j}=b$. Note that we assumed full cross-immunity which is true for a number of variants before the Omicron sub-variants [38]. If there is only a partial cross immunity, $\bar{S}_{i}(t)\neq\bar{S}_{j}(t)$. However, in a short interval, the new immunity (reduction is susceptibility) created by the variants is small. As a result, it can be shown that $\bar{S}_{i}(t)\approx\beta_{ij}^{\prime}\bar{S}_{j}(t)$ for some constant $\beta_{ij}^{\prime}$ which can be absorbed in the parameter $\beta_{ij}$ above. Solving the above recurrence relation, we get:
@@
The parameter $S_{ij}$ can be referred to as the relative growth rate of variant $i$ over variant $j$ [37]. Now, consider a scenario in two regions $A$ and $B$$A$ has the variant $i$ emerging over the previously dominant variant $j$; but variant $j$ never reached region $B$ where variant $k$ is dominant. This makes it difficult to assess the potential impact of variant $i$ on region $B$ when we only know $S_{ij}$. To deal with this challenge, we note that the parameter $S{ij}$ should ideally be independent of the region (this may not hold due to the simplifying assumptions above, so we actually have a region-specific relative growth rate $S^{(p)}_{ij}$ for region $p$). Therefore, we can attempt to find the growth-rate of any variant $i$ with a fixed variant 00, specifically, the original COVID-19 variant. For ease of notation, we represent the growth advantage of variant $i$ with respect variant 00 as $S_{i}$, with $S_{0}=0$, and refer to these values as global growth rates. Ideally, $S_{ij}=S_{i}-S_{j}$, irrespective of the region where the $S_{ij}$ is estimated. To estimate these the global growth rates for all the variants that exist by the time $t$ over all regions $p$, we can set up the following system of linear equations:
@@ III-D1 Graph Construction
The graph construction is illustrated in Figure 3. First, we represent each country as a node in the graph and edges in the graph are established to represent relationships between the countries. There are a total of 87 countries in which at least one variant appears and becomes dominant in the CoVariants dataset [12]. While not shown in Figure 3, self-loops are included to account for internal transmissions. Next, we consider temporal aspect of the problem. The graph may evolve over time to capture changing relationships or influences among countries. This is introduced into our graph as dynamic edge weights based on border control data in the OxCGRT dataset [39]. Section IV-A provides more details on this dataset and its pre-processing.
@@ III-D2 Temporal Encoded Graph Convolutional Network (T-GCN)
In order to extract both temporal and geographical information of the data, we propose a hybrid architecture summarized in Figure 4. In this model, we employ a simple Gated Recurrent Unit (GRU) [40] to extract temporal information of the variant transmissions in $T$ time steps. We temporarily remove the growth rate $S_{i}$, which is not time-dependent, from node features and embed the 1-dimensional (1D) time-dependent features (prevalence ratios) into a 2D latent feature vector. Note that in this configuration, we embed the features separately for each country, treating the nodes as distinct input samples grouped into a single batch, resulting in a 3D latent feature vector. The 3D latent feature vector is then flattened to obtain a final 2D embedding, which serves as input for a one-layer Graph Convolutional Network (GCN). The Leaky ReLU activation function introduces non-linearity after the GCN layer. Additionally, two GraphNorm[41] layers are applied after the GRU and GCN layers for normalization, and dropout is incorporated after the GCN layer for regularization. Finally, the embedded features are fed into a Fully Connected (FC) layer to predict delay and emergence for each country.
@@ III-D3 Dynamics-Informed Loss GCN (DIL-GCN)
where N is the number of countries, $\hat{y}_{i}$ is our models output, and $y_{i}$ is the ground truth. We modify this loss function to steer our GCN towards dynamic predictions inspired by the dynamics-based model. This technique is typically used in PINNs [11]. The adjusted loss function is given by:
@@ III-E2 Emergence Prediction
For the problem of emergence prediction, the chosen loss function is the Binary Cross Entropy (BCE) Loss. In addition to this, we note that when a new variant appears, it only reaches a maximum of 38 countries out of the 87, resulting in a large class imbalance. To account for the class imbalances, class weights are applied to the loss function, providing a balanced learning approach. These weights were calculated using the Inverse Class Frequency method [42].
@@ IV-A1 Covariant Data
CoVariants [12] is our primary dataset which provides the current state of variants and mutations of interest for SARS-CoV-2. This data is enabled by GISAID [43] and ranges from August 2020 through October 2023, with a bi-weekly resolution for 36 variants starting from 20A.EU1 to 23F.Omicron, which is current as of this work. CoVariants follows the NextStrain Clade schema [44], in which variants can descend from other variants.
@@ IV-A2 Border Restrictions Data
We considered the border restrictions data provided in the OxCGRT [39] dataset, with details on what level of border controls were enacted by governments in real-time from Jan 1, 2020, to Jan 8, 2023. OxCGRT measured the variations in government responses using their COVID-19 Government Response Stringency Index, which is a simple additive score of nine indicators ranging from school closures to vaccination policies. Of these indicators, we picked international travel closure controls since they best represent the goal of this paper in capturing the emergence and cross-border transmission of variants. These values are encoded in an ordinal scale as follows: 0 - no measures, 1 - screening, 2 - quarantine on high risk region, 3 - ban on high risk regions, 4 - total border closure. The recorded values were then scaled down to a range between 0.1 and 1, with 1 indicating no border restrictions and 0.1 representing complete border closure. Note that 0.1 was chosen over 0, acknowledging the possibility of some mobility between connecting countries due to necessary trade or border crossings via open land routes. Furthermore, given that our CoVariants data extends beyond January 2023, we extend OxCGRT by making the assumption that all border restrictions have remained unchanged since Jan 8, 2023.
@@ IV-A3 Country Graphs
We considered two graphs representing the interconnection between countries. This dataset depicts country neighborhoods according to the Correlates of War (CoW) database. Countries were also deemed neighbors if they share a land or river border or are located within 24 miles of each other across bodies of water. Adjacency: We used the country adjacency dataset [45], which offers a list of countries and their neighboring countries. OpenFlights: This is a collection of regular flights along with their routes mapped from airports around the world [46]. Given the scope of the current work, we primarily focus on airports and the flight route data, and was mainly used during ablation studies. Airports data is a UTF-8 encoded collection of over 14,000 airports, train stations and ferry terminals across the world as of January 2017. It consists of information like city/country, IATA codes, geographic coordinates and the type of the airport (air terminal, ferry station, etc.). Routes data provides information about 67663 directional routes between 3321 airports on 548 airlines. This dataset was used in an ablation study to validate the effectiveness of utilizing these flight routes as edges in our graph creation process.

### v2

@@ I Introduction
The COVID-19 pandemic presented an unprecedented global health crisis, severely affecting millions of people worldwide and demanding swift and effective responses from governments and healthcare providers [WHO]. As the pandemic progressed, multiple variants of COVID-19 emerged, each possessing unique genetic mutations that can significantly impact transmissibility, virulence, and even vaccine efficacy. During much of the COVID-19 epidemic, the surge in cases and severe outcomes (hospitalizations and deaths) have been driven by the emergence of new variants [Wiemken2023]. Consequently, monitoring the emergence and spread of these variants is crucial for devising appropriate public health measures and optimizing containment strategies [WHO2]. A critical factor driving the surge in a given region is the time of arrival of the new variant [Markov2023]. We can observe how a new variant, that has not appeared in region $A$ yet, spreads in region $B$. We can study the spread in the region $B$ to understand the properties of this new variant. If it is a highly transmissible or immune-evading variant [lambrou2022genomic], we can expect it to spread in the region $A$ eventually. However, when it happens in region $A$ precisely remains unknown. Therefore, making a good prediction of arrival time would lead to a more effective preparation and resource management.
@@
It naturally follows from the above description that Problem 1 is a classification problem while Problem 2 is a regression problem. In practical scenarios, we envision addressing Problem 1 first, and only tackling Problem 2 if the outcome is positive. This is also important for the sake of evaluation as we can answer Problem 2 only if the variant actually arrives in the region of interest. Existing work has focused on the problem of predicting the prevalence of a variant when it has already arrived in the region of interest [sun2020forecasting, hu2020early, wu2020nowcasting]. Specifically, logistic regression turns out to provide a good estimate of the prevalence over time [lr01, lr02]. However, these techniques require the variant to have a non-zero proportion already in the region of interest. To the best of our knowledge, no prior work exists predicting the delay to reach a certain prevalence even when there is zero prevalence in the region. We show that Problem 1 (Emergence prediction) is a simpler problem (albeit, non-trivial) that can be solved by a decision tree. However, Problem 2 (delay detection) is more difficult to solve and requires more sophisticated techniques.
When a new variant emerges with evolutionary favorable properties, the only way it can be transferred to a different region is through a host (human). This encourages the idea of using an underlying network of mobility to address the proposed problems. Since reaching a certain prevalence may depend on the other currently circulating variants, their dynamics (how fast one variant can spread over others) also play a role. Therefore, we propose a variant dynamics-informed Graph Neural Network (GNN) that utilizes a network of mobility and features inspired by variant dynamics to solve the proposed problems. This is different than typical Physics-Informed Neural Networks (PINNs) where the dynamics act as a regularization for the loss [raissi2019physics]. Here, we show that our approach of constructing appropriate features results in lower errors than incorporating the dynamics in the loss function.
Our key contributions represent a novel effort in addressing the challenge of predicting the emergence of COVID-19 variants at a global level, and in doing so establishing a new benchmark for evaluating Emergence Detection and Delay Prediction on the CoVariants dataset [covariants]. More specifically, our contributions are as follows:
@@ II-A Variants and Variant Dynamics
In the context of infectious diseases, the term ”variant” refers to a version of a virus with some changes in its genetic material, known as genome [CDC_class]. These changes happen through genetic mutations and can affect the virus’s characteristics, like how easily it spreads, how severe the illness it causes is, and whether it can evade the immune system or not. Variants can be categorized in multiple ways based on genetic differences, and the significance of these differences can vary. Some categorizations focus on a common ancestor, while others may specifically highlight mutations in key regions of the virus’s genome [CDC_class]. In our study, we rely on the categorizations provided by the CoVariants dataset, discussed further in Section IV-A.
@@
Furthermore, it’s worth noting that the proportions of different variants can be analyzed over time. Some studies have observed that these proportions follow a straight line when plotted on a semi-logarithmic scale. Therefore, they leverage this fact to understand the dynamics of variant emergence and prevalence in populations [Beesley2023]. We also observe this in our data as shown in Figure 2. However, to the best of our knowledge, existing studies do not derive any dynamics capable of predicting delays in the emergence of variants between multiple regions/countries.
@@ II-B Graph Neural Networks and Epidemics
In response to the COVID-19 pandemic, many efforts were made to develop forecasting models capable of predicting the progression and future trends of the COVID-19 pandemic [US, EU, DU]. However, these forecasting approaches have primarily focused on predicting overall case counts, hospitalizations, or mortality rates, rather than predicting delays in variant emergence. The ability to predict the timing of variant-specific outbreaks can significantly enhance public health preparedness and response strategies, allowing for timely interventions such as targeted vaccination campaigns and tailored public health measures [WHO2].
Many Deep Learning (DL) techniques have been explored for predicting the dynamics of infectious diseases. While many models use methods like GNNs [scarselli2008graph], and PINNs [pinnraissi2019physics] for forecasting, some approaches have also focused on genetic sequence modeling for variants [Nagpal2022, Basu2021].
E. Kharazmi et al. [pinnkharazmi2021identifiability], use PINNs in epidemiological models to better capture the dynamics through the underlying Ordinary Differential Equations (ODEs). These works are normally variations of the classical Susceptible-Infectious-Removed (SIR) model with additional compartments [pinnkiselev2023delay]. Extending this, A. Rodríguez et al. [pinnrodriguez2023einns] propose a framework for forecasting, Epidemiologically-informed Neural Networks (EINNs) that integrates epidemiological knowledge embedded in ODE models into neural networks. The core idea is utilizing the learned representations from a PINN using transfer learning along with a neural network that ingests multi-modal data. Other extensions include Epi-DNNs from X. Ning et al. [pinnning2023epi], in which the neural network expresses the unknown parameters in the compartmental model, and the ODEs are solved using the Runge-Kutta method.
Since infectious diseases inherently present across multiple regions over a period of time, network models, like GNNs help in capturing the low-level graph effects. Works [GTNN, STAN, transfergraph, Sung2022, pinnganesan2021spatio], including Gao et al. [STAN] rely on GNNs to predict the total number of cases for a fixed period in the future using a spatio-temporal attention network. These models typically map locations (e.g., a county or a state) to nodes on a graph and construct the edges based on geographical proximity and demographic similarity between locations. As an extension, [pinnseo2019physics, pinnseo2019differentiable] combines spatio-temporal GNNs with PINNs. [chopra2022differentiable, chopra2021deepabm] use agent-based models on top of these to further capture low-level interactions between populations in different compartments.
While some works do focus on using the spatio-temporal networks with location aware features [pinndeng2019graph], including mobility data [pinnkapoor2020examining], to the best of our knowledge, no studies have attempted to predict the emergence of a variant in one country based on its occurrence in another (Problems 1 & 2), through variant informed dynamics.
@@ III-B Variant Growth Rates
We first derive the dynamics of two competing variants in one region. While logistic regression has been widely used to fit these dynamics, for completeness, we present a derivation. We start by using a model of infectious diseases that generalizes a large class of models [SIkJalpha]:
@@
The above further assumes that the mean serial intervals of the two variants are close, i.e., $b_{i}=b_{j}=b$. Note that we assumed full cross-immunity which is true for a number of variants before the Omicron sub-variants [Chen2023]. If there is only a partial cross immunity, $\bar{S}_{i}(t)\neq\bar{S}_{j}(t)$. However, in a short interval, the new immunity (reduction is susceptibility) created by the variants is small. As a result, it can be shown that $\bar{S}_{i}(t)\approx\beta_{ij}^{\prime}\bar{S}_{j}(t)$ for some constant $\beta_{ij}^{\prime}$ which can be absorbed in the parameter $\beta_{ij}$ above. Solving the above recurrence relation, we get:
@@
The parameter $S_{ij}$ can be referred to as the relative growth rate of variant $i$ over variant $j$ [SIkJalpha]. Now, consider a scenario in two regions $A$ and $B$$A$ has the variant $i$ emerging over the previously dominant variant $j$; but variant $j$ never reached region $B$ where variant $k$ is dominant. This makes it difficult to assess the potential impact of variant $i$ on region $B$ when we only know $S_{ij}$. To deal with this challenge, we note that the parameter $S{ij}$ should ideally be independent of the region (this may not hold due to the simplifying assumptions above, so we actually have a region-specific relative growth rate $S^{(p)}_{ij}$ for region $p$). Therefore, we can attempt to find the growth-rate of any variant $i$ with a fixed variant 00, specifically, the original COVID-19 variant. For ease of notation, we represent the growth advantage of variant $i$ with respect variant 00 as $S_{i}$, with $S_{0}=0$, and refer to these values as global growth rates. Ideally, $S_{ij}=S_{i}-S_{j}$, irrespective of the region where the $S_{ij}$ is estimated. To estimate these the global growth rates for all the variants that exist by the time $t$ over all regions $p$, we can set up the following system of linear equations:
@@ III-D1 Graph Construction
The graph construction is illustrated in Figure 3. First, we represent each country as a node in the graph and edges in the graph are established to represent relationships between the countries. There are a total of 87 countries in which at least one variant appears and becomes dominant in the CoVariants dataset [covariants]. While not shown in Figure 3, self-loops are included to account for internal transmissions. Next, we consider temporal aspect of the problem. The graph may evolve over time to capture changing relationships or influences among countries. This is introduced into our graph as dynamic edge weights based on border control data in the OxCGRT dataset [hale2021global]. Section IV-A provides more details on this dataset and its pre-processing.
@@ III-D2 Temporal Encoded Graph Convolutional Network (T-GCN)
In order to extract both temporal and geographical information of the data, we propose a hybrid architecture summarized in Figure 4. In this model, we employ a simple Gated Recurrent Unit (GRU) [cho2014learning] to extract temporal information of the variant transmissions in $T$ time steps. We temporarily remove the growth rate $S_{i}$, which is not time-dependent, from node features and embed the 1-dimensional (1D) time-dependent features (prevalence ratios) into a 2D latent feature vector. Note that in this configuration, we embed the features separately for each country, treating the nodes as distinct input samples grouped into a single batch, resulting in a 3D latent feature vector. The 3D latent feature vector is then flattened to obtain a final 2D embedding, which serves as input for a one-layer Graph Convolutional Network (GCN). The Leaky ReLU activation function introduces non-linearity after the GCN layer. Additionally, two GraphNorm[GraphNorm] layers are applied after the GRU and GCN layers for normalization, and dropout is incorporated after the GCN layer for regularization. Finally, the embedded features are fed into a Fully Connected (FC) layer to predict delay and emergence for each country.
@@ III-D3 Dynamics-Informed Loss GCN (DIL-GCN)
where N is the number of countries, $\hat{y}_{i}$ is our models output, and $y_{i}$ is the ground truth. We modify this loss function to steer our GCN towards dynamic predictions inspired by the dynamics-based model. This technique is typically used in PINNs [raissi2019physics]. The adjusted loss function is given by:
@@ III-E2 Emergence Prediction
For the problem of emergence prediction, the chosen loss function is the Binary Cross Entropy (BCE) Loss. In addition to this, we note that when a new variant appears, it only reaches a maximum of 38 countries out of the 87, resulting in a large class imbalance. To account for the class imbalances, class weights are applied to the loss function, providing a balanced learning approach. These weights were calculated using the Inverse Class Frequency method [classBalance].
@@ IV-A1 Covariant Data
CoVariants [covariants] is our primary dataset which provides the current state of variants and mutations of interest for SARS-CoV-2. This data is enabled by GISAID [gisaid] and ranges from August 2020 through October 2023, with a bi-weekly resolution for 36 variants starting from 20A.EU1 to 23F.Omicron, which is current as of this work. CoVariants follows the NextStrain Clade schema [nextstrain-clade], in which variants can descend from other variants.
@@ IV-A2 Border Restrictions Data
We considered the border restrictions data provided in the OxCGRT [hale2021global] dataset, with details on what level of border controls were enacted by governments in real-time from Jan 1, 2020, to Jan 8, 2023. OxCGRT measured the variations in government responses using their COVID-19 Government Response Stringency Index, which is a simple additive score of nine indicators ranging from school closures to vaccination policies. Of these indicators, we picked international travel closure controls since they best represent the goal of this paper in capturing the emergence and cross-border transmission of variants. These values are encoded in an ordinal scale as follows: 0 - no measures, 1 - screening, 2 - quarantine on high risk region, 3 - ban on high risk regions, 4 - total border closure. The recorded values were then scaled down to a range between 0.1 and 1, with 1 indicating no border restrictions and 0.1 representing complete border closure. Note that 0.1 was chosen over 0, acknowledging the possibility of some mobility between connecting countries due to necessary trade or border crossings via open land routes. Furthermore, given that our CoVariants data extends beyond January 2023, we extend OxCGRT by making the assumption that all border restrictions have remained unchanged since Jan 8, 2023.
@@ IV-A3 Country Graphs
We considered two graphs representing the interconnection between countries. This dataset depicts country neighborhoods according to the Correlates of War (CoW) database. Countries were also deemed neighbors if they share a land or river border or are located within 24 miles of each other across bodies of water. Adjacency: We used the country adjacency dataset [countryAdjacencyData], which offers a list of countries and their neighboring countries. OpenFlights: This is a collection of regular flights along with their routes mapped from airports around the world [OpenFlights2023]. Given the scope of the current work, we primarily focus on airports and the flight route data, and was mainly used during ablation studies. Airports data is a UTF-8 encoded collection of over 14,000 airports, train stations and ferry terminals across the world as of January 2017. It consists of information like city/country, IATA codes, geographic coordinates and the type of the airport (air terminal, ferry station, etc.). Routes data provides information about 67663 directional routes between 3321 airports on 548 airlines. This dataset was used in an ablation study to validate the effectiveness of utilizing these flight routes as edges in our graph creation process.